OpenStudy (anonymous):

For the following system of equations what is the x-value? -x+2y=6 6y=x+18 A.-18 B.0 C.-6 D.3

OpenStudy (anonymous):

-x+2y=6 +x +x 2y=6+x divide both sides by 2 y=(1/2)x+3 now plug it in for y in the other equation 6((1/2)x+3)=x+18 3x+18=x+18 -18 -18 3x=x -x -x 2x=0 divide each side by 2 x=0

OpenStudy (anonymous):

Thank you so much!! @sudukogirl13

OpenStudy (anonymous):

ur welcome

OpenStudy (kewlgeek555):

1. Isolate one of the variables into one of the equations: So you have: -x + 2 y= 6 x + 18 = 6y Let's start with x + 18 = 6y. \[x + 18 = 6y\]\[6y-18=x\]So now we know what x equals. Well, not exactly, but enough to know how to solve for y: 2. Substitute the expression for the isolated variable into the other equation and solve for the other variable. So the other equation is:\[-x+2y=6\]\[-(6y-18)+2y=6\]See what I did there? Mmhmm. Oh, I was doing it the hard way and sudoko girl got it. T.T

OpenStudy (anonymous):

Well thank you to you to !;)